### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
Modern university research labs increasingly rely on workplace instant-messaging platforms such as Slack and Microsoft Teams to facilitate communication and collaboration among lab members, including those working remotely (Perkel, 2017; Wang et al., 2022). Slack alone reported over 1.2 million users across more than 3,000 higher education institutions during the COVID-19 pandemic in 2020 (Slack, 2020). These platforms serve as channels through which lab members share knowledge crucial and practical to their work—e.g., academic writing tips, internship opportunities, and travel grant information—on an ongoing basis (Isaacs et al., 2002; Anders, 2016; Wang et al., 2022). However, this knowledge is often lost. Messages are stored chronologically and without semantic structure, making it difficult to retrieve relevant information once it is buried in long conversation threads (Zhang et al., 2017; Zhang and Cranshaw, 2018; Pasad et al., 2020). As a result, lab directors may receive the same questions repeatedly or give the same advice repeatedly, e.g., correcting punctuation in latex. Students are left to learn through trial and error (Anderson et al., 1996).
@@
To address this challenge, we present CHOIR (Chat-based Helper for Organizational Intelligence Repository), a chatbot-based organizational memory management system embedded in Slack. The goal of CHOIR is to give students access to organizational memory within their communication platform, i.e., Slack, and to allow professors to maintain documents within the platform, leveraging communication that they have in the platform to retrieve, extract, and preserve knowledge. To accomplish this, CHOIR integrates four interconnected interactions: Document-grounded Q&A, which retrieves knowledge with relevant references from a collection of documents when a user asks a question; Q&A Sharing for Follow-up Discussion, which allows users to share Q&A with peers to address unanswered questions or extend discussions; Knowledge Extraction from Conversation, which captures new knowledge that naturally emerges from Slack messages; and AI-assisted Document Update, which facilitates document maintenance by helping managers locate relevant documents and suggest edits. In this way, CHOIR establishes an organizational memory practice that leverages lab members’ communication in Slack to retrieve and sustain knowledge, reducing lab directors’ burden of continuously updating documents while providing members with immediate and convenient access to organizational memory. While this approach can be potentially applicable to many different types of organizations, such as an online community and a company, we chose research labs as our target environment because they constantly experience member turnover, rely on regular communication, and maintain various types of documentation that require collaborative updates (Tuma et al., 2021).
@@
We followed a three-stage approach: (1) a formative study with faculty and graduate students (N=15) to understand their organizational memory practice, if any, challenges in transfering and retreiving knowledge amont lab members; (2) the design and implementation of CHOIR guided by these findings; and (3) a one-month field deployment (N=21) across four university research labs, preceded by document collection and structured onboarding. Based on our interaction trace analysis and follow-up interviews after the field study, we confirm that CHOIR effectively helped students obtain answers available in the documents. In addition, it encouraged professors with no prior experience in maintaining organizational memory to document new knowledge that students wanted to learn or needed to know when questions were shared publicly. However, students faced psychological barriers to sharing their questions—either from concerns about bothering their peers and professors, or from reluctance due to how others might perceive them for not knowing the answers to seemingly basic questions. Similar barriers was noted for answering other people’s questions. Based on the findings, we discuss the social and technical aspects of how organizational memory can be effectively integrated into chat communication platforms for collective contributions, while taking into account students’ hesitation and fear of updating organizational memory documents.
@@
    Empirical evidence demonstrating the effectiveness of a chat-based organizational memory system, along with insights from a month-long deployment that reveal real-world challenges and design implications.
@@ 2. Related Work
In this section, we explore prior research on organizational memory and knowledge management systems, examining how digital platforms have evolved to support knowledge preservation and sharing. We then investigate approaches to transforming conversational data into organizational knowledge, and finally review the emerging role of chatbots and large language models (LLMs) in facilitating knowledge work.
@@ 2.1. From Repository Models to Distributed Organizational Memory
With the advancement of web technologies, wiki systems demonstrated this potential through collaborative editing (Wagner, 2004), but organizational deployments encountered familiar social barriers. Danis (Danis and Singer, 2008) deployed ResearchWiki in a research organization but found limited participation due to competitive culture and modification anxiety. Subsequent research identified systemic barriers: additional effort requirements, information sensitivity concerns, discomfort with unfinished work publication, and openness anxiety (Holtzblatt et al., 2010). The collaborative editing model effective in public contexts created stress in organizational environments where authorship and accountability mattered (Grudin and Poole, 2010). Alternative platforms like Google Docs achieved better adoption through simplified user experiences that naturally facilitated role differentiation without wiki complexity (Sun et al., 2014), with supporting technologies focusing on transparent collaboration history to navigate social complexities (Wang et al., 2015).
@@ 2.4. Summary and Connection to CHOIR
This review reveals consistent evolution in OM systems: from repository models to distributed cognitive systems, from formal documentation to conversational capture, and from manual curation to AI-assisted transformation. However, persistent challenges have remained consistent across decades—incentive misalignment, expert burden, participation barriers, and information overload—indicating that technical advances alone are insufficient. CHOIR addresses this gap by synthesizing three research streams: embedding OM in existing social systems, leveraging existing conversational interactions, and exploiting LLMs’ capabilities. By operating within Slack where research teams already communicate, CHOIR avoids traditional adoption barriers while implementing document-grounded Q&A, conversational knowledge extraction, and AI-assisted documentation workflows. The system design incorporates lessons from prior research: mutually reinforcing Q&A and documentation cycles, integration with existing communication flows, and AI assistance that reduces rather than increases user burden. CHOIR thus represents both technological synthesis and an approach to addressing the persistent social challenges that have limited OM system adoption in collaborative research environments.
@@ 3.1. Method
We focused on recruiting professors and university students to understand the needs and challenges of organizational memory. We recruited participants through a screening survey distributed to faculty and graduate students at research-oriented universities. We advertised the study on the author’s social media to recruit professors in other research-oriented universities. We recruited professors who directed research labs with more than 4 Ph.D. students and had experience or interest in maintaining organizational memory for their research lab. We adapted the eligibility criteria for recruiting students accordingly. From the survey responses, we contacted and recruited 14 participants: 11 faculty members (P1-P11) and 4 graduate students (S1-S4) representing various academic disciplines, primarily from Computer Science 7) but also from other disciplines: Industrial Engineering (2), Information Science (1), and Human Development and Family Science (1). Faculty participants had been running their research labs for an average of 6.4 years (SD = 3.78, range: 4-13 years) and supervised an average of 5 Ph.D. students (SD = 1.67, range: 2-7 students). We also recruited 4 graduate students from 4 different disciplines. Some of the participants had a well-established organizational memory practice, whereas others relied on more informal organizational methods. This study was approved by the internal review board of the author’s university([IRB-XXXXX]).
@@ 3.2.4. Documentation as a Collective Effort with Barriers
In both cases, they noted that students’ ownership of the document could be beneficial, both in encouraging them to read it and in fostering a sense of contribution or achievement. Still, professors wanted to track all changes to ensure that edits made by students were legitimate. Ultimately, professors wished to remain aware of the changes and retain the final decision on whether to approve them.
@@
Overall, professors viewed it positively when students developed a sense of ownership and contributed to building organizational memory collectively. Prior work has shown that granting editing privileges and fostering a sense of community ownership can encourage sustained contributions and effective use of shared documents (Wagner, 2004; Forte and Bruckman, 2005; Morgan et al., 2013). At the same time, professors emphasized the importance of tracking all changes and retaining final approval rights to ensure accuracy, reflecting a recurring tension also observed in large-scale collaborative systems. In Wikipedia, for instance, the balance between openness and quality control has been identified as a central challenge  (Geiger and Ribes, 2010; Arazy et al., 2011). Building on these insights, our design aims to encourage lab members’ participation in curating organizational memory while ensuring that professors can maintain oversight and quality control.
@@ 4. CHOIR: Chat-based Helper for Organizational Intelligence Repository
Grounded in our formative study, we designed CHOIR to enable students to obtain answers directly within Slack and to transform everyday conversations into a maintained organizational memory. CHOIR integrates four features essential for intelligent knowledge retrieval and documentation: Document-grounded Q&A, Q&A Sharing for Follow-up Discussion, Knowledge Extraction from Conversation, and AI-assisted Document Update. These features reflect the design goals identified in the formative study, and the underlying implementation utilizes LLM-powered multi-agents that update and access documents stored in GitHub via GitHub and Slack API.
When a message is sent to CHOIR—either through a direct message or by mentioning its handle (i.e., @CHOIR) in a channel any channel that it is part of, the message will be classified as either a question or an update request, corresponding to the two start nodes shown in Figure 2. The four features in CHOIR can establish a feedback loop of how a question asked by a student can eventually update the lab handbook. For example, the member who posed a question may decide to share the question with their peers, and the peers can answer the question, suggesting updating the document. Finally, the manager will decide whether to apply updates. The flow may also begin outside of Q&A, when a member or the manager directly provides new knowledge to CHOIR. CHOIR’s four features will be detailed in the remainder of this section, showing how each aligns with a design goal (DG1–DG4).
@@ 4.1. Document-grounded Q&A with CHOIR
CHOIR allows lab members to ask a question to it and retrieve information from the document in any Slack channel. When a question is submitted, CHOIR answers by composing a response grounded in those sources (DG2). Members can submit questions to CHOIR either by sending a direct message or by mentioning it in a Slack channel.
Internally, CHOIR retrieves topically relevant chunks and performs generation conditioned on the member’s question and the retrieved text. To avoid hallucination, CHOIR includes an answerability guard: when the repository does not provide sufficient coverage, CHOIR abstains from inventing an answer, surfaces related passages with citations, and clarifies what information is available (e.g., related sections or partial guidance). This makes documentation gaps visible for further update, and keeps replies aligned with current documents.
Each Q&A exchange is posted with grounding and provenance so that the questioner can quickly verify sources and decide whether to continue the discussion. Figure 3 highlights four UI elements: the member’s question ( A in Figure 3, submitted via direct message in this example), CHOIR’s answer in the main message ( B in Figure 3), and an evidence panel in the thread that lists the specific chunks used with anchored links ( C in Figure 3). These references encourage the questioner to check out the original documents available in a GitHub repository, thereby getting familiar with the organizational memory.
If a question is asked in a public channel, all members can naturally see the CHOIR’s answers, establishing a common ground that supports further discussion or follow-up questions and thereby facilitates communication among members (DG3).
After CHOIR provides a response, users are given the option to continue the discussion with their peers if they have follow-up questions, regardless of whether the original question was fully answered. A private message that is only visible to the questioner ( D in Figure 3) provide two buttons with which the questioner can choose either to post the Q&A exchange to the designated Q&A channel for lab-wide follow-up or to send it privately to selected peers or the manager via DM, creating a natural handoff to the next feature.
@@ 4.2. Q&A Sharing for Follow-up Discussion
CHOIR encourages questioners to relay a question—whether answered or unanswered— to a Slack channel or to a specific person via a group chat, thereby encouraging communication among members (DG3). After CHOIR posts an answer, the questioner sees a share banner offering two options to share Q&A either in a public channel (Green button in Figure 3- D) or send a direct message (“Ask in Private” D, Figure 3)) to a private group chat with selected members, CHOIR, and the questioner.
Questioners can specify more options in one of two modal windows shown in Figure 4. The Your Comment field ( A in Figure 4) allows the questioner to add context or a follow-up prompt. For the “Ask in Private” option, the question asker can choose lab members that they will share the question with using a People picker ( D in Figure 4), which is pre-populated with the manager by default but can be edited to include any Slack members (e.g., senior Ph.D. students). Privacy Options ( B in Figure 4) enable anonymous sharing, in which case CHOIR will replace the questioner’s name with “A team member” when the message is shared in a Q&A channel to stay anonymous. In addition, when anonymity is enabled in the private-sharing modal, the questioner is also excluded from the group chat conversation to preserve anonymity among recipients. The preview panel ( C in Figure 4) shows exactly how the question will be shared, either as a channel post or as an outgoing DM.
The full Q&A exchange will be shared as shown in the left half of the Figure 5, and it appears as a message that includes the question and CHOIR’s answer, along with the questioner’s comment and profile image ( A in Figure 5). Other members or the manager can then add follow-up comments ( B in Figure 5). The same posting format also applies when the Q&A is shared in a private DM with selected people via the ”Ask in Private” modal. While answers are posted in a public channel or if a questioner is part of the group chat, the questioner can see how others responded to the relayed question. If a message is shared anonymously in a private DM, CHOIR forwards any follow-up replies to the original questioner without revealing their identity. CHOIR is designed to encourage Q&A sharing so that it supplements rather than replaces interpersonal communication in the lab, enabling members to collectively answer questions. More importantly, when such discussions lead to new or missing knowledge, Q&A sharing can provide the entry point for the next feature, Knowledge Extraction from Conversation.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
Modern university research labs increasingly rely on workplace instant-messaging platforms such as Slack and Microsoft Teams to facilitate communication and collaboration among lab members, including those working remotely (Perkel, 2017; Wang et al., 2022). Slack alone reported over 1.2 million users across more than 3,000 higher education institutions during the COVID-19 pandemic in 2020 (Slack, 2020). These platforms serve as channels through which lab members share knowledge crucial and practical to their work—e.g., academic writing tips, internship opportunities, and travel grant information—on an ongoing basis (Isaacs et al., 2002; Anders, 2016; Wang et al., 2022). However, this knowledge is often lost. Messages are stored chronologically and without semantic structure, making it difficult to retrieve relevant information once it is buried in long conversation threads (Zhang et al., 2017; Zhang and Cranshaw, 2018; Pasad et al., 2020; Meixner et al., 2017). As a result, lab directors may receive the same questions repeatedly or give the same advice repeatedly, e.g., correcting punctuation in latex. Students are left to learn through trial and error (Anderson et al., 1996).
@@
To address this challenge, we present CHOIR (Chat-based Helper for Organizational Intelligence Repository), a chatbot-based organizational memory management system embedded in Slack. The primary goal of CHOIR is twofold: (1) to build and sustain a research lab’s organizational memory, and (2) to enable lab members to easily access knowledge from it. To that end, our approach embeds organizational memory directly within their communication platform—Slack—allowing students to ask questions to a chatbot and allowing professors to modify documents within the platform, leveraging knowledge that emerges from their conversations. To accomplish these goals, CHOIR integrates four interconnected interactions: Document-grounded Q&A, which retrieves knowledge with relevant references from a collection of documents when a user asks a question; Q&A Sharing for Follow-up Discussion, which allows users to share Q&A with peers to address unanswered questions or extend discussions; Knowledge Extraction from Conversation, which captures new knowledge that naturally emerges from Slack messages; and AI-assisted Document Update, which facilitates document maintenance by helping managers locate relevant documents and suggest edits. In this way, CHOIR supports an organizational memory practice that leverages lab members’ communication in Slack to retrieve and sustain knowledge, reducing lab directors’ burden of continuously updating documents while providing members with immediate and convenient access to organizational memory. While this approach can be potentially applicable to many different types of organizations, such as an online community and a company, we chose research labs as our target environment because they constantly experience member turnover, rely on regular communication, and maintain various types of documentation that require collaborative updates (Tuma et al., 2021).
@@
We followed a three-stage approach: (1) a formative study with faculty and graduate students (N=15) to understand their organizational memory practice, if any, challenges in transferring and retrieving knowledge among lab members; (2) the design and implementation of CHOIR guided by these findings; and (3) a one-month field deployment (N=21) across four university research labs, preceded by document collection and structured onboarding. Based on our interaction trace analysis and follow-up interviews after the field study, we found that CHOIR helped students obtain answers available in the documents. In addition, it supported professors with no prior experience in maintaining organizational memory to document new knowledge that students wanted to learn or needed to know when questions were shared publicly. However, our findings revealed a fundamental tension between students’ privacy preferences and directors’ need for awareness of knowledge gaps. Students faced psychological barriers to sharing their questions—either from concerns about bothering their peers and professors, or from reluctance due to how others might perceive them for not knowing the answers to seemingly basic questions. Similar barriers were noted for answering other people’s questions and for contributing to documentation updates. Based on the findings, we discuss the design implications for AI-mediated organizational memory systems, including the privacy-awareness trade-off, the role of AI as a facilitator for human communication, and approaches to expand the perceived boundary of documentable knowledge.
@@
    Findings from a month-long deployment that reveal real-world usage patterns, challenges and design implications for AI-mediated organizational memory systems.
@@ 2. Related Work
We explore prior research on organizational memory and knowledge management systems, examining how digital platforms have evolved to support knowledge preservation and sharing. We then investigate approaches to transforming conversational data into organizational knowledge, and finally review the emerging role of chatbots and large language models (LLMs) in facilitating knowledge work.
@@ 2.1. From Repository Models to Distributed Organizational Memory
With the advancement of web technologies, wiki systems demonstrated this potential through collaborative editing (Wagner, 2004), but organizational deployments encountered familiar social barriers. Danis (Danis and Singer, 2008) deployed ResearchWiki in a research organization but found limited participation due to competitive culture and modification anxiety. Similarly, Munson (Munson, 2008) found in a MediaWiki deployment that while the system motivated documentation of past projects and enabled discovery of organizational work, members felt uncomfortable making substantial edits to others’ content. Broader studies of enterprise wikis found that making workers participate in wiki is challenging and organization’s culture that support collaboration is necessary (Kiniti and Standing, 2013). Subsequent research identified systemic barriers: additional effort requirements, information sensitivity concerns, discomfort with unfinished work publication, and openness anxiety (Holtzblatt et al., 2010). The collaborative editing model effective in public contexts created stress in organizational environments where authorship and accountability mattered (Grudin and Poole, 2010). Alternative platforms like Google Docs achieved better adoption through simplified user experiences that naturally facilitated role differentiation without wiki complexity (Sun et al., 2014), with supporting technologies focusing on transparent collaboration history to navigate social complexities (Wang et al., 2015).
@@ 2.4. Summary and Connection to CHOIR
This review reveals consistent evolution in OM systems: from repository models to distributed cognitive systems, from formal documentation to conversational capture, and from manual curation to AI-assisted transformation. However, persistent challenges have remained consistent across decades—incentive misalignment, expert burden, participation barriers, and information overload—indicating that technical advances alone are insufficient. Prior AI-based systems such as Tilda and Collabot primarily address subprocesses of the OM lifecycle, mainly capturing and summarizing conversation (Zhang and Cranshaw, 2018; Tepper et al., 2018), but do not support the full cycle: how extracted knowledge is validated, integrated into OM documents, and easily retrieved by the members. CHOIR addresses this gap by synthesizing three research streams: embedding OM in existing social systems, leveraging existing conversational interactions, and exploiting LLMs’ capabilities. By operating within Slack where research teams already communicate, CHOIR avoids traditional adoption barriers while implementing document-grounded Q&A, conversational knowledge extraction, and AI-assisted documentation workflows. The system design incorporates lessons from prior research: mutually reinforcing Q&A and documentation cycles, integration with existing communication flows, and AI assistance that reduces rather than increases user burden. CHOIR thus represents both technological synthesis and an approach to addressing the persistent social challenges that have limited OM system adoption in collaborative research environments.
@@ 3.1. Method
We recruited participants through a screening survey distributed to faculty and graduate students at research-oriented universities. We advertised the study on the author’s social media to recruit professors in other research-oriented universities. We recruited 15 participants: 11 faculty members (P1-P11) from various academic disciplines, primarily Computer Science (7) but also Industrial Engineering (2), Information Science (1), and Human Development and Family Science (1), and 4 graduate students (S1-S4) from different disciplines. We targeted professors who directed research labs with more than 4 Ph.D. students and had experience or interest in maintaining organizational memory. Faculty participants had been running their research labs for an average of 6.4 years (SD = 3.78, range: 4-13 years) and supervised an average of 5 Ph.D. students (SD = 1.67, range: 2-7 students). Some of the participants had a well-established organizational memory practice, whereas others relied on more informal organizational methods. This study was approved by our institution’s IRB([IRB-XXXXX]).
@@ 3.2.1. Challenges in Creating and Maintaining Lab Documentation
“There was one time there was like a whole step just completely missing, and I couldn’t get it to work. […] The people who had written the entry into the Notion had already graduated.” (S4)
@@ 3.2.2. Need for Effective Knowledge Retreival
Communication platforms’ features, such as search function or pinned messages, provide a way to reuse existing messages to some extent. However, they reported difficulties using those features for knowledge retrieval and management.
“I probably don’t use those [search functions] as effectively as I could. So, there’s definitely been times when I’m trying to find an old conversation and I’m searching within a DM from maybe one student and it’s not coming up, and I’m not really sure how I can find it.” (P9)
“Pinning a message[…] it can get quite a large, like, a long list of items. And like the pinned message, you can only see, like, a couple.” (P1)
While existing functions allow users to retrieve or curate past messages they authored, this approach does not scale as the communication history grows, and retrieving information through pinned messages is similarly unsustainable.
@@ 3.2.4. Documentation as a Collective Effort with Barriers
In both cases, they noted that students’ ownership of the document could be beneficial, both in encouraging them to read it and in fostering a sense of contribution or achievement. However, students expressed reluctance toward documentation work, viewing it as tedious.
“The documentation part is always tedious. Nobody wants to do that. […] There are times when you feel like, okay, this could be automated or this could be a dedicated job for an AI agent or something like that.” (S2)
Still, professors wanted to track all changes to ensure that edits made by students were legitimate. Ultimately, professors wished to remain aware of the changes and retain the final decision on whether to approve them.
@@
Overall, professors viewed it positively when students developed a sense of ownership and contributed to building organizational memory collectively. However, students expressed reluctance toward such documentation work, highlighting a tension between professors’ expectations and students’ actual willingness to participate.
Prior work has shown that granting editing privileges and fostering a sense of community ownership can encourage sustained contributions and effective use of shared documents (Wagner, 2004; Forte and Bruckman, 2005; Morgan et al., 2013). At the same time, professors emphasized the importance of tracking all changes and retaining final approval rights to ensure accuracy, reflecting a recurring tension also observed in large-scale collaborative systems. In Wikipedia, for instance, the balance between openness and quality control has been identified as a central challenge  (Geiger and Ribes, 2010; Arazy et al., 2011). Building on these insights, our design aims to encourage lab members’ participation in curating organizational memory while ensuring that professors can maintain oversight and quality control.
@@ 4. CHOIR: Chat-based Helper for Organizational Intelligence Repository
Grounded in our formative study, we designed CHOIR to support students to obtain answers directly within Slack and to transform everyday conversations into a maintained organizational memory. CHOIR’s primary objective of sustaining the growth of—and easy access to—organizational memory translates into several subgoals derived from our formative study. These subgoals include streamlining document updates (DG1), enabling easy and accurate information retrieval (DG2), facilitating communication among members (DG3), and supporting their participation in managing organizational memory (DG4). We accomplish these goals by embedding CHOIR as a chatbot within the lab’s communication platform, i.e., Slack. CHOIR integrates four features for the subgoals: Document-grounded Q&A, Q&A Sharing for Follow-up Discussion, Knowledge Extraction from Conversation, and AI-assisted Document Update. These features reflect the design goals identified in the formative study.
When a message is sent to CHOIR—either through a direct message or by mentioning its handle (i.e., @CHOIR) in any channel that it is part of, CHOIR processes it as either a question or an update request, corresponding to the two start nodes shown in Figure 2. The four features in CHOIR can support a feedback loop of how a question asked by a student can eventually update the lab handbook. For example, the member who posed a question may decide to share the question with their peers, and the peers can answer the question, suggesting updating the document. Finally, the manager will decide whether to apply updates. The flow may also begin outside of Q&A, when a member or the manager directly provides new knowledge to CHOIR. CHOIR’s four features will be detailed in the remainder of this section, showing how each aligns with a design goal (DG1–DG4).
@@ 4.1. Document-grounded Q&A with CHOIR
CHOIR allows lab members to ask a question and retrieve information from the documentation in any Slack channel. When a question is submitted, CHOIR answers by composing a response grounded in those sources (DG2). To avoid hallucination, CHOIR includes an answerability guard: when the repository does not provide sufficient coverage, CHOIR abstains from inventing an answer, surfaces related passages with citations, and clarifies what information is available (e.g., related sections or partial guidance). This makes documentation gaps visible for further update, and keeps replies aligned with current documents.
Each Q&A exchange is posted with grounding and provenance so that the questioner can quickly verify sources. Figure 3 highlights the UI elements: the member’s question ( A), CHOIR’s answer ( B), an evidence panel listing the specific chunks used with anchored links ( C), and a share banner ( D) for further discussion. These references encourage the questioner to check out the original documents, thereby getting familiar with the organizational memory.
@@ 4.2. Q&A Sharing for Follow-up Discussion
After CHOIR provides an answer, the questioner can share the Q&A exchange with others to facilitate communication among members (DG3). The share banner ( D in Figure 3) offers two options: posting to a public Q&A channel or sending via direct message to selected members. Questioners can configure additional options through one of two modal windows shown in Figure 4. The Your Comment field ( A in Figure 4) allows the questioner to add context or a follow-up prompt.
For the “Ask in Private” option, the question asker can choose lab members that they will share the question with using a People picker ( D in Figure 4), which is pre-populated with the manager by default but can be edited to include any Slack members (e.g., senior Ph.D. students). Privacy Options ( B) enable anonymous sharing, in which case CHOIR will replace the questioner’s name with “A team member” when the message is shared in a Q&A channel to stay anonymous. In addition, when anonymity is enabled in the private-sharing modal, the questioner is also excluded from the group chat conversation to preserve anonymity among recipients. The preview panel ( C) shows exactly how the question will be shared, either as a channel post or as an outgoing DM.
The shared Q&A exchange includes the question, CHOIR’s answer, the questioner’s comment, and profile image as shown in Figure 5 ( A), where other members can add follow-up comments ( B). When shared anonymously in a private DM, CHOIR forwards replies to the original questioner without revealing their identity. When such discussions lead to new knowledge, they provide the entry point for Knowledge Extraction from Conversation.